OpenStudy (kirbykirby):

Integral: arctanx/(1+x^2) dx

OpenStudy (kirbykirby):

I get (arctanx)^2/2 but wolfram gives something else

OpenStudy (anonymous):

\[\int\limits \frac{\tan^{-1}x}{1+x^2}dx\]just to dbl check. This the right problem?

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

yeah that's what I am getting as well

OpenStudy (anonymous):

u-substitute for arctanx

OpenStudy (kirbykirby):

Oops.. i just realized I put d/dx instead of the integral sign on wolfram -_- I'm tired :(

OpenStudy (anonymous):

@kirbykirby do u wanto integrate or differentiate

OpenStudy (anonymous):

lmao... noticed that

OpenStudy (anonymous):

integrate

OpenStudy (anonymous):

@@kirbykirby u have integrated it correctly

OpenStudy (anonymous):

so do the u sub

OpenStudy (anonymous):

\[\frac{d}{dx} arctanx=\frac{1}{1+x^2}\]which is eactly what we have there!!

OpenStudy (kirbykirby):

Yeah thanks everyone :)!! I wish I could give a medal to everyone

OpenStudy (kirbykirby):

I'm so silly lol

OpenStudy (anonymous):

hey tyhanks guys @kirbykirby, I'm assuming you posted this in response to my orignal post?

OpenStudy (anonymous):

or we're in the same class lol

OpenStudy (kirbykirby):

What's your original post?

OpenStudy (kirbykirby):

@egr2b I don't know if I know you o_O?

OpenStudy (kirbykirby):

Oh I didn't even know you posted that haha!

OpenStudy (kirbykirby):

Maybe we are in the same class ._. this is weird loool

OpenStudy (anonymous):

lol... good luck then